OpenStudy (msmr):

3sqrt{32} - 4sqrt{63}

OpenStudy (anonymous):

32=2^5 so sqrt{32} = 2^(5/2) 63=9*7= 7*3^2 so sqrt{63} = 3*sqrt{7}

OpenStudy (msmr):

wait...what?

OpenStudy (msmr):

32=2^5?

OpenStudy (anonymous):

\[2^{5}=32\]

OpenStudy (msmr):

okay, but i'm trying to get the square roots to be the same so I can subtract the terms.

OpenStudy (msmr):

I think.

OpenStudy (anonymous):

3*sqrt{32} = 3*2^(5/2) =12sqrt{2} 4*sqrt{63} = 12sqrt{7}

OpenStudy (msmr):

okay thank you...

OpenStudy (anonymous):

u r welcome.